Use the following information: A nurse needs to prepare strength Enfamil for several infants in the nursery. The supply is 8-ounce cans of Enfamil. How much sterile water should be mixed with each 8-ounce can of Enfamil in order to prepare the + strength solution?
Answer #1

Solution: According to the given question the given strength is 1/4 and is clearly stating that

1/4 strength = 1 part Enfamil to 4 parts of solution

Therefore 4 – 1 = 3 (4 parts solution -1 part Enfamil = 3 parts sterile water)

3 parts sterile water = 3 x 8 = 24 ounces of sterile water